# Joint Pricing and Resource Allocation: An Optimal Online-Learning Approach

### v1

@@ Appendix A Proof Details
According to the definition of $W(p)$ given by Eq. 6, we have
@@
Due to the property that the limit operation preserves convexity, it is sufficient to show that $\check{Q}_{\Delta}(\vec{I}^{*},p)$ is convex with respect to $p$. According to Lemma A.6, we know that $\check{Q}_{\Delta}(\vec{I}^{*},p)$ has at most $E:=(\frac{2C}{\Delta})^{n}\cdot 3^{m+n}$ non-smooth singularities. Without loss of generality, denote them as
@@
Proof of Lemma A.8.
We firstly consider $W^{\prime}(P_{e}^{-})$. According to the proof of Lemma A.7, we have
@@
Applying Lemma A.7 on Eq. 24, we know that $W(p)$ is convex within each smooth interval $(P_{e},P_{e+1})$. Also, from Lemma A.8, we know that $W(p)$ is convex at any singularity $P_{e}$ as its left derivatives does not exceed its right derivatives. Combining those two properties, we know that $Q(\vec{I}^{*},p)$ is convex w.r.t. $p$. This ends the proof.
A.5 Proof of Lemma 5.2
@@
Please kindly find the proof of Lemma A.9 in Section A.8.
@@
We will specify the value of $\hat{\epsilon}$ as a function of $\epsilon$ by the end of this proof. We defer the proof of Lemma A.10 to Section A.9. From Lemma A.10, we may get the following corollary.
Corollary A.11.
@@
Proof of Corollary A.11.
For each fixed tuple $(K,\tau,s,\hat{p}_{\tau})$, according to Lemma A.10, we have
@@
Combining Corollary A.11 with Lemma A.9, we have the following corollary
Corollary A.12.
Define a $flag$ as shown in Algorithm 5, and define $p_{K}^{*}:=\mathop{\mathrm{argmin}}_{p\in[C_{i_{K},j_{K}},c_{i_{K+1},j_{K+1}}]}W(p).$ When $flag==0$ by the end of Sub-Epoch $s$ of Epoch $\tau$, we have
@@
Therefore we know that $|\hat{Q}_{K,\tau,s,b_{K,\tau}}-\hat{Q}_{K,\tau,s,c_{K,\tau}}|\leq 4\Delta_{K,\tau,s}$. Combining Corollary A.11, we have
@@
By applying Lemma A.9 with $\Delta=\frac{4C_{c}}{\sqrt{n_{s}}}$, we show that the lemma holds with $\Pr\geq 1-24T\hat{\epsilon}$ (since we have used Corollary A.11 for 6 times). ∎
@@
    Here the fourth line comes from Corollary A.11 and the last line is an application of $C_{c}=\delta_{K}$ and Corollary A.12.

### v2

@@ Appendix B Proof Details
Denote $\vec{I}^{*}(p):=\mathop{\mathrm{argmin}}_{\vec{I}}Q(\vec{I},p)$. According to the definition of $W(p)$ given by Eq. 5, we have
@@
Due to the property that the limit operation preserves convexity, it is sufficient to show that $\check{Q}_{\Delta}(\vec{I}^{*},p)$ is convex with respect to $p$. Also, since $g(\vec{I},p,\vec{D})$ is the value to a linear programming(LP), the solution should be located on the vertex of its feasible space. Therefore, the non-smooth singularities of $g(\vec{I}^{*}(p),p,\vec{a}-\vec{b}\cdot p+\vec{N}_{t})$ only exists when multiple vertices of the LP feasible space coincide. According to Eq. 2, the feasible space have at most ($mn+m+n\choose mn$ vertices. Denote $\mathcal{G}(m,n):=2^{{mn+m+n}\choose mn}$, and we know that there are at most $\mathcal{G}(m,n)$ chances that at least two vertices can coincide with each other (hence causing a non-smooth singularity). As a consequence, we have $\check{Q}_{\Delta}(\vec{I}^{*},p)$ has at most $E:=(\frac{2C}{\Delta})^{n}\cdot\mathcal{G}(m,n)$ non-smooth singularities. Without loss of generality, denote them as
@@
Proof of Lemma B.4.
We firstly consider $W^{\prime}(P_{e}^{-})$. According to the proof of Lemma B.3, we have
@@
Applying Lemma B.3 on Eq. 22, we know that $W(p)$ is convex within each smooth interval $(P_{e},P_{e+1})$. Also, from Lemma B.4, we know that $W(p)$ is convex at any singularity $P_{e}$ as its left derivatives does not exceed its right derivatives. Combining those two properties, we know that $Q(\vec{I}^{*},p)$ is convex w.r.t. $p$. This ends the proof.
B.4 Proof of Lemma 5.2
@@
Please kindly find the proof of Lemma B.5 in Section B.7.
@@
We will specify the value of $\hat{\epsilon}$ as a function of $\epsilon$ by the end of this proof. We defer the proof of Lemma B.6 to Section B.8. From Lemma B.6, we may get the following corollary.
Corollary B.7.
@@
Proof of Corollary B.7.
For each fixed tuple $(K,\tau,s,\hat{p}_{\tau})$, according to Lemma B.6, we have
@@
Combining Corollary B.7 with Lemma B.5, we have the following corollary
Corollary B.8.
Define a $flag$ as shown in Algorithm 3, and define
$p_{K}^{*}:=\mathop{\mathrm{argmin}}_{p\in[C_{i_{K},j_{K}},c_{i_{K+1},j_{K+1}}]}W(p).$ (33)
When $flag==0$ by the end of Sub-Epoch $s$ of Epoch $\tau$, we have
@@
Therefore we know that $|\hat{Q}_{K,\tau,s,b_{K,\tau}}-\hat{Q}_{K,\tau,s,c_{K,\tau}}|\leq 4\Delta_{K,\tau,s}$. Combining Corollary B.7, we have
@@
By applying Lemma B.5 with $\Delta=\frac{4C_{c}}{\sqrt{n_{s}}}$, we show that the lemma holds with $\Pr\geq 1-24T\hat{\epsilon}$ (since we have used Corollary B.7 for 6 times). ∎
@@
    Here the fourth line comes from Corollary B.7 and the last line is an application of $C_{c}=\delta_{K}$ and Corollary B.8.